Development Partners Housing has been practicing cash management for some time by using the Baumol model for determining cash balances. Some time ago, the model called for an average balance (C*/2) of $500; at that time, the rate on marketable securities was 4.0 percent. A rapid increase in interest rates has driven the interest rate up to 9.0 percent. What is the appropriate average cash balance now? Note that the fixed cost per transaction and the total cash needed for transactions during the year did not change. Enter your answer rounded to two decimal places. Do not enter $ or comma in the answer box. For example, if your answer is $12,300.456 then enter as 12300.46 in the answer box.
Answer #1

C*(Old) = $500*2 = $1000

C*(Old) = [2FT / r]1/2

$1,000 = [2FT / 0.04]1/2

2FT / 0.04 = [1,000]2

2FT = 1,000,000 * 0.04 = $40,000

C*(New) = [2FT / r]1/2

= [$40,000 / 0.09]1/2

= [$444,444.44]1/2

= $666.66

Average Cash Balance = C*/2 = $666.66/2 = $333.33